OpenStudy (anonymous):

1. Evaluate the integral: ( 4 Marks) a) b)

OpenStudy (anonymous):

\[a) \int\limits_{0}^{\pi/4} \cos ^{2} \left( 4t-\pi/4 \right)dt\]

OpenStudy (anonymous):

\[b ) \int\limits_{0}^{\pi/2} 3\sin x \cos x \div \sqrt{1+3\sin ^{2}} x\]

OpenStudy (zarkon):

for (a) use the 1/2 angle formula (then substitution) for (b) use substitution \[u=1+3\sin^2(x)\]